Take 95 South, Inc. has the following returns of the last five years, -18.3%, 11.9%, 9.2%, 26.4%, -11.1%. With this information construct the 95% confidence interval for Take 95 South, Inc. returns.
Answer #1

Ans:- For 95% confidence interval the lower limit rage = Mean - Z Value * Standard deviation (SD) and For Upper Limit range = Mean + Z Value * Standard deviation (SD). we will use the AVERAGE and STDEV function of excel to find the Mean and Standard deviation.

Z value for 95% is 1.96 (approx).

K15 fa A B с 1 2 3 Years 4 5 1 2 3 4 Returns (%) -18.30 11.90 9.20 26.40 -11.10 6 7 8 5 9 10 Mean 3.62 11 12 SD 18.14 13 14 LK15 fo A B C 1 2 3 Years Returns (%) -18.3 4 1 5 2. 11.9 6 3 9.2 7 4 26.4 - 11.1 8 5 9 10 Mean =AVERAGE(C4:08) SD ESTDEV(C4:0

Therefore 95% confidence interval for Take 95 South, Inc. returns. is between -31.93% and 39.17%.